A salesman sold $300 more than the rest of the sales staff combined. If the sales total for the day was $2500, then how much did the rest of the sales staff sell?


x+300=2500
x=2200

Answer by nerdybill(5407)

A salesman sold $300 more than the rest of the sales staff combined. If the sales total for the day was $2500, then how much did the rest of the sales staff sell?
.
Let x = amount "rest of sales staff sold"
then
x+300 = amount "salesman sold"
.
"rest of sales staff sold" + "salesman sold" = total sales
x + (x+300) = 2500
2x + 300 = 2500
2x = 2200
x = $1100